OpenStudy (anonymous):

PLEASE HELPPP Determine the type of boundary line and shading for the graph of the inequality -5x - y<= -2 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

terenzreignz (terenzreignz):

If it's <= (there's the -or equal to- extension) you can rule out the dashed lines :) Next, try inputing the origin into your inequality (0,0) -5x - y <= -2 -5(0) - 0 <= -2 0 <= -2 Which is not true So the origin must not be in the shaded part :D So it's a solid line, with shading on the side that does NOT include the origin cheers ;)

OpenStudy (anonymous):

THANKS CAN U HELP ME WITH THIS . "Which of the following is not an equivalent form of the compound inequality x + 18<= 11 and x + 18 =>3 A number line with a closed circle on -15, shading to the left, and a closed circle on -7, shading to the right. 11 => x + 18 =>3 -7 =>x => -15 A number line with a closed circle on -15, a closed circle on -7, and shading in between.

terenzreignz (terenzreignz):

Sorry for the delay, trying to find the best way to explain...

terenzreignz (terenzreignz):

OK x + 18 <= 11 then x <= 11 - 18 x <= -7 and also: x + 18 >= 3 then x >= 3-18 x >= -15 to combine these inequalities, write: -15 <= x <= -7 does this give you a better picture? :)

OpenStudy (anonymous):

so wait the answer is -15 <= x <= -7